# lemmasection

### v1

@@ 6. Evaluation Metrics
Another important observation pertains to the performance metrics used. While macro AUROC remains the standard metric in ECG benchmarks (Strodthoff et al., 2021), it is relatively insensitive to class imbalance, which is a common characteristic of clinical ECG datasets. In contrast, macro AUPRC is more sensitive to this imbalance, as it captures the precision-recall trade-off (Ozenne et al., 2015). This distinction is evident in most of our experiments. For instance, in the four PTB-XL multi-label tasks examined in this study, the proposed method achieves improvements of 1.2%–3.3% in macro AUROC compared to the baseline, whereas the improvement in macro AUPRC ranges from 5.3% to 20.9%. On the one hand, the substantial gains in macro AUPRC highlight the efficacy of the post-training strategy. On the other hand, the notable discrepancy between macro AUROC and AUPRC underscores an important insight: even a small difference in macro AUROC can correspond to significant variations in macro AUPRC, indicating that different methods may exhibit considerable differences in their ability to detect true positives in the presence of a large number of negatives.
7. Implications and Future Directions
@@ 6 Conclusion
In this study, we propose a simple yet effective post-training strategy for the recently released ECG foundation model, ECGFounder. Experiments on the PTB-XL benchmark demonstrate that the proposed method outperforms both the baseline fine-tuning strategy and several recent state-of-the-art models. Key components, including stochastic depth and preview linear probing, were identified as critical drivers of performance improvement. Additionally, the proposed method exhibits greater sample efficiency and training stability. Our findings underscore the potential of post-training strategies to bridge the adaptation gap for ECG foundation models, highlighting a promising direction to further advance their development.

### v2

@@ A General Post-training Framework
Another important observation pertains to the performance metrics used. While macro AUROC remains the standard metric in ECG benchmarks (Strodthoff et al., 2021), it is relatively insensitive to class imbalance, which is a common characteristic of clinical ECG datasets. In contrast, macro AUPRC is more sensitive to this imbalance, as it captures the precision-recall trade-off (Ozenne et al., 2015). This distinction is evident in most of our experiments. For instance, in the four PTB-XL multi-label tasks examined in this study, the proposed method achieves improvements of 0.7%-8.9% in macro AUROC compared to the baseline, whereas the improvement in macro AUPRC ranges from 23.3% to 77.9%. On the one hand, the substantial gains in macro AUPRC highlight the efficacy of the post-training strategy. On the other hand, the notable discrepancy between macro AUROC and AUPRC underscores an important insight: even a small difference in macro AUROC can correspond to significant variations in macro AUPRC, indicating that different methods may exhibit considerable differences in their ability to detect true positives in the presence of a large number of negatives.
Implications and Future Directions
@@ 6 Conclusion
In this study, we propose a simple yet effective post-training strategy for general ECG foundation models. Using a Transformer-based foundation model as an example, we demonstrate the effectiveness of the proposed approach. Experiments on the PTB-XL dataset show that the proposed method not only significantly outperforms the baseline fine-tuning but also maintains stable convergence and strong data efficiency. Moreover, the proposed method achieves competitive or superior performance compared with several recent state-of-the-art models across diverse ECG classification tasks. These findings underscore the potential of post-training strategies to bridge the adaptation gap for ECG foundation models, highlighting a promising direction to further advance their development.